OpenStudy (anonymous):

1/t + 1/p = 1/d solution for t??

OpenStudy (anonymous):

saweet man, just curious im assuming you play music?

OpenStudy (anonymous):

i play guitar a little. but thats it

OpenStudy (anonymous):

nice man, im a music major, definitely not a math! lol

OpenStudy (anonymous):

haha cool. what do u play, or do you sing?

OpenStudy (anonymous):

well im a clarinet and sax major, but do keys and bass on the side,

OpenStudy (anonymous):

damn. thats awesome.

OpenStudy (anonymous):

\[t=-\frac{d p}{d-p} \]

OpenStudy (anonymous):

\[\frac{1}{t} = \frac{1}{d} - \frac{1}{p}\]\[\implies \frac{1}{t} = \frac{p}{dp} - \frac{d}{dp}\]\[\implies \frac{1}{t} = {p - d \over dp} \]\[\implies t = {dp \over p-d}\] Though what rob said is also the same thing.. Just more 'simplified' without the extra negative.

OpenStudy (anonymous):

Yes, I agree. I will try to conform next time. Thank you for your observation.

OpenStudy (anonymous):

@tricketts8737 Don't you think you could spare a medal for polpak? You seem to be passing them out for invalid or non related answers.